# General Seemingly Unrelated Local Projections

### v2

@@ 4. Step 4:
In our empirical illustrations, we cycle through our algorithm $12,000$ times and discard the first $3,000$ draws as burn-in. All inference is then based on each $3$rd of the remaining draws. Given the conditionally conjugate structure of our model, mixing for most parameters appears to be no issue in most cases, as measured by inefficiency factors. The only exceptions are the hyperparameters associated with the kernel of the GP. In this case, mixing is slightly worse but still acceptable. For our most general specification with multiple shocks, it takes about $14$ seconds to obtain $1,000$ draws from the posterior on a Macbook Air M1.
@@ 4. Empirical Applications
In all of our empirical work below, we do not consider contemporaneous control variables, i.e., $\bm{r}_{t}=\emptyset$. The LPs, which Equation (7) specifies in levels, are estimated using long differences, see, e.g., Piger and Stockwell (2023):
@@
The implementation with a “flat” prior sets $\bm{\mu}_{\beta}=\bm{0}_{H}$ and $\bm{V}_{\beta}=10\bm{I}_{H}$. The default LP is estimated with OLS using Newey and West (1987) HAC standard errors. The smooth LP follows Barnichon and Brownlees (2019), with a penalty matrix that shrinks towards a line ($r=2$ in their notation). Consistent with the original paper, we select the hyperparameters using cross-validation.
4.1. Simulated data
We simulate $T=250$ observations from the Smets and Wouters (2003) dynamic stochastic general equilibrium model, and estimate the response of output to a monetary policy shock. The structural parameters of the model are obtained along the lines of Smets and Wouters (2003) and implemented in the R package gEcon (see Klima et al., 2015).
(a) SU-LP
(b) SU-LP (flat)
(c) LP (default)
(d) LP (smooth)
Figure 3: Comparison of estimated dynamic effects across different LP implementations. The grey shaded area marks the 68/90/95 percent credible set alongside the median (solid black line), and analogous confidence intervals and mean estimates for classical versions. The dashed red lines are the posterior median of the GP prior alongside a 90 percent credible set. The green line is the true impulse response function.
Figure 3 plots the output responses for our benchmark case with a GP prior, a flat prior in the impulse response coefficients, OLS with HAC standard errors, as well as the smoothed LPs of Barnichon and Brownlees (2019). For our approach, we also plot the posterior for the Gaussian process in red, which we can interpret as an estimate of a smoothed LP.
[…]
Figure 4 plots the estimated (median) posterior covariance matrix of the forecast errors along with its Cholesky factor. We show these results to highlight that indeed there seems to be a pattern in this matrix that is broadly consistent with the results from our earlier examples. More generally, our approach captures the comovement between forecast errors across horizons.
4.2. Monetary policy in the US
@@
Figure 5: Estimated monetary policy shocks, their time-varying log-variance, and resulting relevance statistic. The grey shaded area marks the 68/90/95 percent credible set alongside the median (solid black line) and observed instrument (blue circles).
@@
Figure 6: Estimated dynamic effects of a monetary policy shock on real GDP for SU-LP estimated with a GP prior and NG shrinkage. The grey shaded area marks the 68/90/95 percent credible set alongside the median (solid black line). The dashed red line is the posterior median of the GP prior.
@@
Figure 7: Comparison of estimated dynamic effects across different LP implementations. The grey shaded area marks the 68/90/95 percent credible set alongside the median (solid black line) for Bayesian implementations (and analogous confidence intervals and mean estimates for classical implementations). The dashed red lines are the posterior median of the GP prior alongside a 90 percent credible set.
@@ 4.3. Monetary and fiscal policy in the US
Figure 8: Time series and scatter plots of the standardized instruments, monetary policy (MP) and two variants for fiscal policy (FP); “cor” is the correlation coefficient.
@@
Figure 9: Comparison of estimated dynamic effects when assuming correlated measurement errors of monetary and fiscal shock instruments. The grey shaded area marks the 68/90/95 percent credible set alongside the median (solid black line). The dashed red lines are the posterior median of the GP alongside a 90 percent credible set.
@@
Figure 10: Comparing the response of real GDP to the monetary shock under different model specifications. Solid lines are the posterior median alongside the 68% credible set, and the dashed lines are the smooth GP estimate. The top row shows the single-shock baseline specification (estimated with GP and NG prior, shocks with iid measurement error, and stochastic volatility) in gray. The blue variants by panel vary the indicated complication holding everything else fixed. The lower panels show a comparison to the joint-shock (including the fiscal shock of Fisher and Peters, 2010) case (left panel), and correlated versus independent measurement errors of the instruments (right panel).
@@
Figure 11: Estimating responses to the joint component of the monetary policy and fiscal shocks. “PC” extracts the first principle component from the instruments and includes it as a single instrument; “Common factor” estimates the joint component in our unified framework.
@@ 5. Conclusions
We have presented a general framework for inference in LPs that exploits advantages of the Bayesian toolkit. Our framework allows for multiple mismeasured shocks, can estimate a common shock that partially determines commonly used instruments in macroeconomics, allows for flexible, yet parsimonious regularization, requires minimal tuning of prior hyperparameters, directly models the correlation structure present in multi-step forecast errors, and automatically takes into account missing data. Finally, estimation is fast, removing one of the commonly cited advantages of standard frequentist inference in LPs.

### v3

@@ 4. Step 4:
In our empirical illustrations, we cycle through our algorithm $12,000$ times and discard the first $3,000$ draws as burn-in. All inference is then based on each 33rd of the remaining draws. Given the conditionally conjugate structure of our model, mixing for most parameters appears to be no issue in most cases, as measured by inefficiency factors. The only exceptions are the hyperparameters associated with the kernel of the GP. In this case, mixing is slightly worse but still acceptable. For our most general specification with multiple shocks, it takes about $14$ seconds to obtain $1,000$ draws from the posterior on a Macbook Air M1.
3.7. Pseudo likelihoods, mis-specification and power posteriors
Using a pseudo likelihood $\hat{p}(\bm{y}|\bm{\Xi})$, with $\bm{\Xi}$ denoting all model parameters (and eventual static factors integrated out analytically), instead of the true (but unknown) likelihood function implies misspecification if the two differ significantly from each other. As discussed in the previous sub-sections, SU-LP offers two mechanisms which aim to control for misspecification. These are the system-based estimation approach that controls for past forecast errors and the non-parametric prior on $\bm{\beta}$ that provides additional flexibility. We will show in the next section that these features indeed lead to empirical coverage rates that are close to the nominal ones.
However, if the researcher believes that this is not enough (for instance, if the DGP features structural breaks, non-linear features, or non-Gaussian shocks) one option would be to robustify SU-LP towards different (but unknown) forms of misspecification. A solution that is popular in the literature on Bayesian statistics builds on using the so-called power posterior (Holmes and Walker, 2017; Grünwald and Van Ommen, 2017):
$p_{c}(\bm{\Xi}|\bm{y})=\frac{\hat{p}(\bm{y}|\bm{\Xi})^{c}~p(\bm{\Xi})}{p(\bm{y}|c)},$ (20)
for a learning rate $c\in(0,1]$. A value of 0 implies that no learning takes place (and the power posterior equals the prior) while $c=1$ gives the standard posterior obtained through the MCMC algorithm outlined in Section 3.6. Intermediate values imply that the likelihood is downweighted and less weight is put on data-based information. Essentially, smaller values of $c$ protect posterior inference against significant levels of misspecification whereas if we believe that the model is well specified, setting $c$ close to one are appropriate.
In simple models, using the power posterior is easy and posterior simulation can be carried out conditional on a particular value of $c$. However, in our hierarchical model that consists of multiple layers (with several latent quantities), we opt for a different approach that takes the standard posterior (i.e., with $c=1$) and ex-post modifies them to end up with draws from the posterior for $c<1$. Our approach builds on importance sampling and uses the standard posterior as a importance density and weights that depend on $\hat{p}(\bm{y}|\bm{\Xi})^{c-1}$. [^17] For a similar approach to setting the learning rate of a power posterior, see Sona et al. (2023).
These weights are obtained by dividing Equation (20) by $p(\bm{\Xi}|\bm{y})=\hat{p}(\bm{y}|\bm{\Xi})p(\bm{\Xi})/p(\bm{y})$, noting that the ratio of the power marginal likelihood $p(\bm{y}|c)$ and the standard marginal likelihood $p(\bm{y})$ does not depend on $\bm{\Xi}$, leads to:
$\frac{p_{c}(\bm{\Xi}|\bm{y})}{p(\bm{\Xi}|\bm{y})}\propto\hat{p}(\bm{y}|\bm{\Xi})^{c-1}.$
If we’d like to compute $E_{c}(g(\bm{\Xi}))=\int g(\bm{\Xi})p_{c}(\bm{\Xi}|\bm{y})d\bm{\Xi}$, multiplying by $\hat{p}(\bm{\Xi}|\bm{y})\times 1/\hat{p}(\bm{\Xi}|\bm{y})$ results in:
$E_{c}(g(\bm{\Xi}))$ $=\int\frac{g(\bm{\Xi})p_{c}(\bm{\Xi}|\bm{y})p(\bm{\Xi}|\bm{y})}{p(\bm{\Xi}|\bm{y})}d\bm{\Xi},$
which, after using $\hat{p}(\bm{y}|\bm{\Xi})^{c}=\hat{p}(\bm{y}|\bm{\Xi})^{c-1}\times\hat{p}(\bm{y}|\bm{\Xi})$, equals
$E_{c}(g(\bm{\Xi}))=\int g(\bm{\Xi})\hat{p}(\bm{y}|\bm{\Xi})^{c-1}p(\bm{\Xi}|\bm{y})d\bm{\Xi},$ (21)
which is the importance sampling identity with the proposal distribution $p(\bm{\Xi}|\bm{y})$ and $p_{c}(\bm{\Xi}|\bm{y})$ being the target density.
Let $\bm{\Xi}^{(s)}$ denote the $s^{th}$ draw from the SU-LP posterior. Given a sequence of $S$ draws, we can approximate (21) as:
$\hat{E}_{c}(g(\bm{\Xi}))=\sum_{s=1}^{S}g(\bm{\Xi}^{(s)})~w^{(s)},$
where $w^{(s)}=\hat{p}(\bm{y}|\bm{\Xi}=\bm{\Xi}^{(s)})^{c-1}/\sum_{j}\hat{p}(\bm{y}|\bm{\Xi}=\bm{\Xi}^{(j)})^{c-1})$. After obtaining weights, the posterior $p_{c}(\bm{\Xi}|\bm{y})$ can then be approximated numerically by sampling (with replacement) from $p(\bm{\Xi}|\bm{y})$ with weights $\bm{w}=(w_{1},\dots,w_{S})^{\prime}$. Adding this step involves minimal computational overhead. However, as is common for importance sampling techniques, it could be that the weights become degenerate. This implies that the effective sample size of the power posterior is small and hence a much larger number of draws from the standard posteriors need to be sampled. But given that this needs to be done once we consider it as relatively unproblematic.
There are several ways to choose the value of $c$. Some methods — such as the SafeBayes approach of Grünwald and Van Ommen (2017) — require cross-validation, which can make the selection of $c$ computationally intensive. Fortunately, in our setting, we can evaluate $p_{c}(\bm{\Xi}\mid\bm{y})$ easily for a range of values of $c$, allowing for alternative strategies. One option is to choose $c$ such that the posterior credible intervals of $p_{c}(\bm{\beta}\mid\bm{y})$ match the width of uncertainty bands produced by commonly used robust LP estimators. Alternatively, we may select $c$ to ensure that the resulting credible intervals are conservative while still yielding statistically significant impulse responses at the horizons of interest. In that case, the chosen value of $c$ is itself informative, quantifying the amount of evidence required to detect a significant effect. A third approach is to use simulations based on a realistic data-generating process and select $c$ to minimize the distance between actual and nominal coverage rates.
4. A Monte Carlo study
4.1. Monte Carlo design
In this section, we analyze the performance of SU-LP (and the variant based on the power posterior) using a realistic DGP that aims to mimic the dynamics of US macroeconomic data. Following Schorfheide (2005) and González-Casasús and Schorfheide (2025), we assume a VARMA($P,\infty$) process for an $n$-dimensional vector of observables $\bm{w}_{t}$, for a sample $t=1,\ldots,T$:
$\bm{w}_{t}=\sum_{p=1}^{P}\bm{\Phi}_{p}\bm{w}_{t-p}+\bm{H}\bm{\epsilon}_{t}+\alpha T^{-\pi}\sum_{j=1}^{\infty}\bm{A}_{j}\bm{H}\bm{\epsilon}_{t-j},\quad\bm{\epsilon}_{t}\overset{\text{iid}}{\sim}\mathcal{N}(\bm{0},\bm{I}_{n}),$
with $n\times n$ dynamic coefficients $\bm{\Phi}_{p}$, $\bm{H}$ is an $n\times n$ structural impact matrix, and $\bm{A}_{j}$ are $n\times n$ MA coefficients. The weight on the MA part is a function of the sample size, implying that the importance of the MA term declines with increasing sample sizes. The term $\alpha$ measures the overall weight; if the researcher estimates a standard VAR, then $\alpha$ measures the overall degree of misspecification and $\pi$ measures local misspecification.
To closely mimic US macroeconomic dynamics, we partially calibrate the parameters using a reduced form VAR(P)(P). We use $n=7$ macroeconomic and financial variables transformed towards approximate stationarity for the US economy, choose $P=5$ lags, and obtain $\bm{H}$ with a Cholesky decomposition of the reduced-form covariance matrix. Subsequently, we use the posterior median estimate of any calibrated parameters, set $\pi=0.5$, and the MA coefficients are simulated from independent standard normal distributions (up to a maximum order of $10$, for $j>10$ they are set equal to 0).
We rely on $1000$ replications for $T+1000$ periods and discard the first $1000$ observations, which leads to samples of length $T$. We consider four sample sizes, $T\in\{100,250,500,1000\}$, and two cases for the weight on the MA part, $\alpha\in\{0,2\}$. We compute impulse responses up to a maximum horizon $\tilde{H}=16$, and refer to the true IRF as $\bm{\beta}_{\ast}$ below.
Similarly to our applications with real data, we focus on the responses of the output growth variable to a shock in the monetary policy rate. Moreover, we consider the output responses for four models. The first is the proposed GP and GL prior setup, referred to as SU-LP. The second is an implementation with a “flat” prior that sets $\bm{\mu}_{\beta}=\bm{0}_{H}$ and $\bm{V}_{\beta}=10\bm{I}_{H}$, labeled SU-LP (flat). The third is the classical LP estimated with OLS using Newey and West (1987) HAC standard errors (LP, default). The fourth is the regularized classical LP, following Barnichon and Brownlees (2019), with a penalty matrix that shrinks towards a line ($r=2$ in their notation; labeled LP, smooth). Consistent with the original paper, we select the hyperparameters using cross-validation.
Additional details such as specific details on the DGP calibration, the computation of true IRFs, and associated charts of example DGP realizations are provided in Appendix A. Appendix B includes another small-scale simulation exercise based on a DSGE-based DGP.
4.2. Monte Carlo results
Table 2 shows the coverage probabilities of the 90 percent posterior credible set or confidence interval across different LP implementations for $1000$ replications of the DGP for selected horizons. To avoid mixing up effects of structural identification, measurement errors, and statistical estimation, we treat the true shock as observed for all competing specifications in this analysis.
The table suggests that for very short samples ($T=100$) and $\alpha\in\{0,2\}$, both variants of SU-LP produce coverage rates that are much closer to the nominal 90 percent level than the OLS-based LPs with HAC standard errors and the smooth LPs of Barnichon and Brownlees (2019), both of which produce coverage rates substantially below 90 percent. Comparing both SU-LP specifications reveals that the specification with a flat prior on $\bm{\beta}$ performs slightly better for horizons up to six steps ahead. For larger impulse response horizons, the pattern is mixed and shrinking the IRFs towards a Gaussian process produces slightly more favorable coverage ratios for $h\in\{8,10,12\}$ before producing slightly too wide credible intervals for longer-run IRFs.
Table 2: Coverage probabilities of the 90 percent posterior credible set or confidence interval across different LP implementations for $1000$ replications of the DGP for selected horizons. Sample size $T\in\{100,250,500,1000\}$ and weight on MA part $\alpha\in\{0,2\}$. Gray shading indicates undercoverage, red shading vice versa.
When we increase the length of the sample to $T=250$ and $T=500$, LP (default) produces better calibrated credible intervals with coverage rates closer to 90 percent. However, both variants of SU-LP yield superior coverage rates for (most) forecast horizons considered. Interestingly, we find no discernible differences between DGPs that set $\alpha=0$ or $\alpha=2$.
The comparatively strong performance of SU-LP also carries over to very large samples ($T=1000$). In this case, SU-LP and SU-LP (flat) produce coverage rates close to 90 percent. LP (default) also produces well calibrated intervals. Only LP (smooth) undercovers and produces too narrow credible intervals, a finding that is consistent with simulation results in Barnichon and Brownlees (2019).
Coverage rates tell us whether our model produces credible intervals surrounding the LPs that are well calibrated. They tell us nothing about the error we make when producing point estimates of the IRFs. In principle, a successful approach should produce a low bias under relatively low risk (defined as the standard deviation of the estimator). We investigate this relationship systematically in Figure 3. The figure shows the median (and 25th and 75th percentiles for SU-LP variants) absolute bias, defined as $\mathbb{E}(\bm{\beta}-\bm{\beta}_{\ast})$, and the standard deviation $\text{Var}(\bm{\beta})^{1/2}$ normalized by $(\bm{\beta}_{\ast}^{\prime}\bm{\beta}_{\ast}/\tilde{H})^{1/2}$, inspired by Li et al. (2024), for the different LP variants across $1000$ realizations from the DGP.
Starting with the bias in the top panel of Figure 3, there are four notable insights. First, regardless of the value of $\alpha$ and all sample sizes, there are only minor differences in bias with respect to impact estimates. However, when we extend the impulse response horizon, a consistent picture emerges where SU-LP produces the lowest bias for $h\geq 4$. Second, when only bias is considered, the default LP estimator exhibits the weakest performance among all models considered. In all cases and for most horizons, it is the approach that produces the largest bias. Third, comparing SU-LP and SU-LP (flat) shows that forcing the IRF estimator towards a Gaussian process pays off, particularly for longer impulse response horizons. LP (smooth) displays a performance comparable to that of SU-LP (flat). Fourth, relative differences across DGPs are small and we mostly find that if the DGP suggests misspecification, most models produce a larger bias. However, the main exception is SU-LP which, irrespective of the value of $\alpha$, produces biases that are barely distinguishable from each other for different values of $\alpha$ (but conditional on $T$). This corroborates our narrative that adding a GP component to the model alleviates bias arising from misspecification.
Li et al. (2024) state that bias reduction is no free lunch, meaning that methods that produce low bias do so at the cost of more estimation uncertainty. Intuitively speaking, we would expect methods that perform well in terms of bias to be accompanied by larger standard deviations. However, this only holds for SU-LP (flat) and (with an opposite sign) for the standard LP estimator. Particularly for SU-LP, we find that is produces a low bias with the second lowest standard deviation across all DGPs considered and for (most) horizons considered. This, in combination with the coverage rates, paints a favorable picture of the statistical properties of SU-LP which produces accurate IRF estimates with reasonably calibrated credible intervals.
Fig. 3: Median (and 25th and 75th percentile for SU-LP variants) absolute bias $\mathbb{E}(\bm{\beta}-\bm{\beta}_{\ast})$ and standard deviation $\text{Var}(\bm{\beta})^{1/2}$ normalized by $(\bm{\beta}_{\ast}^{\prime}\bm{\beta}_{\ast}/\tilde{H})^{1/2}$ across $1000$ replications of the DGP. Sample size $T\in\{100,250,500,1000\}$ and $\alpha\in\{0,2\}$.
4.3. Assessing power posteriors
Our Monte Carlo results suggest that SU-LP performs well across common types of misspecification considered in the literature. However, there may be other types of misspecification that are even more severe. In this case, the researcher can use the power posterior outlined in Section 3.7.
In this section, we evaluate how the power posterior performs for various $c$ values. Our emphasis remains on coverage rates, bias, and standard deviations, particularly examining the SU-LP configuration with GP and GL priors. The Monte Carlo analysis indicated negligible (qualitative) differences in SU-LP’s performance across different sample sizes, so we use $T=250$, which is similar to the number of observations available in common quarterly macroeconomic datasets.
Our results across $1000$ draws from the DGP and for $\alpha\in\{0,2\}$ are structured similar to the preceding simulation study. Figure 4 shows coverage rates of the 90 percent posterior credible set in panel (a) for different settings of the coarsening parameter $c\in\{0.80,0.81,\ldots,1.00\}$. Note that $c=1$ refers to the standard posterior. In panel (b) we plot the normalized median absolute bias and standard deviation for selected values of $c$.
Panel (a) of Figure 4 suggests that if the DGP features no MA term, using the power posterior hurts in terms of coverage as long as $c<0.85$. For $c=0.85$, we obtain coverage rates close to 90 percent for most forecast horizons. Notice that for $h\leq 4$, the standard posterior produces coverage rates very close to 90 percent as well. Only in the case of longer-run IRFs ($h\geq 8$) we find that the model based on the power posterior produces coverage rates close to the nominal one, whereas $c=1$ produces slightly inflated credible intervals. This story, albeit to a lesser degree, carries over to the DGP that sets $\alpha=2$. In this case, we again find that $\alpha=0.85$ is a good choice in terms of achieving coverage close to $90$ percent. The standard posterior performs well throughout, with coverage rates close to (or slightly) above $90$ percent.
Fig. 4: Coverage probabilities of the 90 percent posterior credible set in panel (a) for different settings of the coarsening parameter $c\in\{0.80,0.81,\ldots,1.00\}$. Median absolute bias $\mathbb{E}(\bm{\beta}-\bm{\beta}_{\ast})$ and standard deviation $\text{Var}(\bm{\beta})^{1/2}$ normalized by $(\bm{\beta}_{\ast}^{\prime}\bm{\beta}_{\ast}/\tilde{H})^{1/2}$ in panel (b). Results across $1000$ replications with sample size $T=250$ and $\alpha\in\{0,2\}$.
Considering bias and standard deviation (see panel (b)) reveals that the standard posterior is highly competitive, producing a bias close to the one of the power posterior with $c=0.85$ but with substantially smaller standard deviations. This holds for both values of $\alpha$.
In summary, our simulation results demonstrate that SU-LP maintains robust performance, even in DGPs with MA terms. While introducing the power posterior might offer some improvements, the increase in coverage is generally minor. As for bias and standard deviations, such enhancements are even less significant. Consequently, when using US macroeconomic time series data similar to the one we use to calibrate the DGP, we suggest using the standard posterior.
@@ 5. Empirical Applications
In all of our empirical work below, we do not consider contemporaneous control variables, i.e., $\bm{r}_{t}=\varnothing$. We produce IRFs using the same four LP estimators (SU-LP, SU-LP (flat), LP (default) and LP (smooth)) as in the Monte Carlo exercise.
5.1. Monetary policy in the US
In our empirical work with real data, the LPs, which Equation (7) specifies in levels, are estimated using long differences, see, e.g., Piger and Stockwell (2023):
@@ 5.1. Monetary policy in the US
Fig. 5: Estimated monetary policy shocks, their time-varying log-variance, and resulting relevance statistic. The gray shaded area marks the 68/90/95 percent credible set alongside the median (solid black line) and observed instrument (blue circles).
@@
Fig. 6: Estimated dynamic effects of a monetary policy shock on real GDP for SU-LP estimated with a GP prior and NG shrinkage. The gray shaded area marks the 68/90/95 percent credible set alongside the median (solid black line). The dashed red line is the posterior median of the GP prior.
@@
Fig. 7: Comparison of estimated dynamic effects across different LP implementations. The gray shaded area marks the 68/90/95 percent credible set alongside the median (solid black line) for Bayesian implementations (and analogous confidence intervals and mean estimates for classical implementations). The dashed red lines are the posterior median of the GP prior alongside a 90 percent credible set.
@@ 5.2. Monetary and fiscal policy in the US
Fig. 8: Time series and scatter plots of the standardized instruments, monetary policy (MP) and two variants for fiscal policy (FP); “cor” is the correlation coefficient.
@@
Fig. 9: Comparison of estimated dynamic effects when assuming correlated measurement errors of monetary and fiscal shock instruments. The gray shaded area marks the 68/90/95 percent credible set alongside the median (solid black line). The dashed red lines are the posterior median of the GP alongside a 90 percent credible set.
@@
Fig. 10: Comparing the response of real GDP to the monetary shock under different model specifications. Solid lines are the posterior median alongside the 68% credible set, and the dashed lines are the smooth GP estimate. The top row shows the single-shock baseline specification (estimated with GP and NG prior, shocks with iid measurement error, and stochastic volatility) in gray. The blue variants by panel vary the indicated complication holding everything else fixed. The lower panels show a comparison to the joint-shock (including the fiscal shock of Fisher and Peters, 2010) case (left panel), and correlated versus independent measurement errors of the instruments (right panel).
@@
Fig. 11: Estimating responses to the joint component of the monetary policy and fiscal shocks. “PC” extracts the first principle component from the instruments and includes it as a single instrument; “Common factor” estimates the joint component in our unified framework.
@@ 6. Conclusions
We have presented a general framework for inference in LPs that exploits advantages of the Bayesian toolkit. Our framework allows for multiple mismeasured shocks, can estimate a common shock that partially determines commonly used instruments in macroeconomics, allows for flexible, yet parsimonious regularization, requires minimal tuning of prior hyperparameters, directly models the correlation structure present in multi-step forecast errors, and automatically takes into account missing data. Relative to existing work, our approach delivers superior coverage probabilities in realistic environments. Finally, estimation is fast, removing one of the commonly cited advantages of standard frequentist inference in LPs.